Please help!!!
(4x+4)/5 = (5x-1)/2 - x/3 You must show all work to receive full credit.
Show work here:






Answer by checkley71(8403)

(4X+4)/5=(5X-1)/2-X/3 FIND A COMMON DENOMINATOR FOR THE RIGHT HAND TERMS.
(4X+4)/5=[3(5X-1)-2X]/2*3
(4X+4)/5=(15X-3-2X)/6
(4X+4)/5=(13X-3)/6 NOW CROSS MULTIPLY.
6(4X+4)=5(13X-3)
24X+24=65X-15
24X-65X=-15-24
-41X=-39
X=-39/-41
X=39/41 OR .95 NSWER.